What is the molecular mass of a non-electrolyte solute in a solution that contains 0.6kg of solute and

200g of benzene with a freezing point equal to 5.4°C? The freezing point of pure benzene is 5.53°C/m.

The freezing point constant of benzene is 5.12°C.


1
Expert's answer
2022-03-30T17:35:03-0400

"\Delta T=(5.53-5.4)\degree C=0.13\degree C"

"\Delta T=C_m\cdot K_f=(m/M)/m_s\cdot K_f=>"

"M=m/m_s\cdot K_f/\Delta T=600/200\cdot 5.12/0.13=118.15 g/mol"